I am having trouble with word problems. The questions is: Separate 180 into two parts so that one part will be 3 1/2 times the other. I am not sure what to do, can you help me? Thank you!
Found 2 solutions by lwsshak3, nerdybill:
Answer by lwsshak3(11628)

I am having trouble with word problems. The questions is: Separate 180 into two parts so that one part will be 3 1/2 times the other. I am not sure what to do, can you help me? Thank you
**
let x = one part
Other part = 3 1/2*x=7x/2
x+7x/2=180
2x+7x=360
9x=360
x=40
other part=7*40/2=140

Answer by nerdybill(7384)

Separate 180 into two parts so that one part will be 3 1/2 times the other.
.
Let x = other part
then from "one part will be 3 1/2 times the other"
3.5x = one part
.
x + 3.5x = 180
4.5x = 180
x = 180/4.5
x = 40
.
"one part":
3.5x = 140
.
answer: 40 and 140